Session: Feb 2006 (session 1)
Wavelength: 5 cm


To give feedback on this experiment, click

Experiment Feedback

Onsala (Michael Lindqvist at Sat Feb 25 13:03:41 2006) Success :
RFI seen in the bandpass around 13.45 UTC. Otherwise no known problems.

Medicina (Sergio Mariotti at Sat Feb 25 13:38:02 2006) Minor failures : LATE ANT
Observation started 1h30' later due to Antenna technical problems

Torun (Kaz Borkowski at Mon Feb 27 06:27:35 2006) Success :
No known problems.

Westerbork (Tony Foley at Mon Feb 27 12:50:04 2006) Success :
Single dish was used. Please change the coordinates in the vexfile accordingly. No known problems

Effelsberg (A. Polatidis at Tue Feb 28 10:29:46 2006) Success :
Missed scan 1150-1201 UT for pointing

Jodrell_Bank (Alastair Gunn at Fri Mar 3 17:14:55 2006) Failure :
The MkII telescope was used (Jb2). Disk pack JOD-0023 subsequently developed "directory flawed or missing" error (during experiment ea035a). No data visible on disk pack but may be recoverable. Telescope off source between 20:01 and 20:13 UT due to azimuth wrap limits. Telescope parked at 03:46 for remainder of experiment due to high winds.

Cambridge (Alastair Gunn at Fri Mar 3 17:15:19 2006) Success :
No known problems.

Noto (P. Cassaro at Tue Mar 7 15:53:36 2006) Success :

EVN experiment feedback facility: entire session

Entire Session

Session: Feb 2006 (session 1)
Wavelengths: 6 3.6 5 18 cm


To give feedback on this entire session, click

Entire Session Feedback

Torun (Kaz Borkowski at Mon Feb 20 11:25:29 2006) Minor failures : PCAL
Tr phase cal system failed to produce signals during all the 6 cm experiments for yet unknown reason.

Westerbork (Tony Foley at Fri Feb 24 15:48:24 2006) Minor failures : ANT POIN
The feedback from the antenna control for SINGLE-DISH observations is unreliable for 20seconds after a source change. We will try to fix this. (Affected only the 5cm experiments -Z.P.)